OpenStudy (johnnydicamillo):

The length of a rectangle is 1m more than double the width, and the area of the rectangle is 66m^2. Find the demensions of the rectangle. Please explain the process rather than just giving an answer.

OpenStudy (anonymous):

No it's yours

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

 \[X \times ( 2X + 1) = 66\] \[2X ^{2} +X- 66 = 0 \] and get the value of positive x as the length can't be negative.